OpenStudy (anonymous):

MORE MEDALZ! Solve. |x| + 3 = 7 A. {–10} B. {4} C. {–4, 4} D. {–10, –4}

OpenStudy (anonymous):

|x| + 3 = 7 Subtract 3 to both sides, what's 7 - 3?

OpenStudy (r_josh10):

B. 4 because you subtract 7 from 3.

OpenStudy (anonymous):

7-3=4

OpenStudy (anonymous):

@r_josh10 Don't give answers, we have to have the asker's participation..plus that's wrong.

OpenStudy (anonymous):

@killit!gentlythough. Yes, so our answer is -4 and 4, because the absolute value sign will make -4 positive.

OpenStudy (texaschic101):

absolute values are always positive, even when they say they are negative

OpenStudy (anonymous):

Unless the negative sign is outside of the absolute value sign.

OpenStudy (anonymous):

{x} means how far the number is from the number line.

OpenStudy (anonymous):

okay thanks you guys!

OpenStudy (texaschic101):

YW :)

OpenStudy (r_josh10):

Actually A negativeplus a positive is a ngative

OpenStudy (anonymous):

No problem. :)

OpenStudy (anonymous):

What does that have to do with this question? @r_josh10

OpenStudy (anonymous):

one moar qustin

OpenStudy (r_josh10):

what

OpenStudy (anonymous):

Select the graph that represents the inequality. 2x + 3y < 6

OpenStudy (r_josh10):

add negative 3 to both side

OpenStudy (r_josh10):

then divide by both side. and theres your answer

OpenStudy (r_josh10):

Let's solve for x. 2x+3y<6 Step 1: Add -3y to both sides. 2x+3y+−3y<6+−3y 2x<−3y+6 Step 2: Divide both sides by 2. 2x 2 < −3y+6 2 x< −3 2 y+3 Answer: x< −3 2 y+3

OpenStudy (anonymous):

graphs are here:

OpenStudy (r_josh10):

Theres only on graph.

OpenStudy (texaschic101):

2x + 3y < 6 temporarily put = in for < 2x + 3y < 6 3y < -2x + 6 y < -2/3x + 2 slope is -2/3 y intercept is (0,2) x intercept can be found by subbing in 0 for y 2x + 3y = 6 2x + 3(0) = 6 2x = 6 x = 3....x intercept is (3,0) so the intercepts are (3,0) and (0,2).....and since it is less then with no equal signs, you use open circles with shading below the line

OpenStudy (texaschic101):

nvm what I said about temp removing the < sign

OpenStudy (anonymous):

line is dashed right?

OpenStudy (texaschic101):

line is dashed because there is no equal sign..correct

OpenStudy (anonymous):

I got it. thank you!

OpenStudy (texaschic101):

the graph you posted looks correct except it needs a dashed line